# Topological consequences of null-geodesic refocusing and applications to $Z^{x}$ manifolds

### v2

@@ Abstract
Let $(M,h)$ be a connected, complete Riemannian manifold, let $x\in M$ and $l>0$. Then $M$ is called a $Z^{x}$ manifold if all geodesics starting at $x$ return to $x$ and it is called a $Y^{x}_{l}$ manifold if every unit-speed geodesic starting at $x$ returns to $x$ at time $l$. It is unknown whether there are $Z^{x}$ manifolds that are not $Y^{x}_{l}$-manifolds for some $l>0$. By the Bérard-Bergery theorem, any $Y^{x}_{l}$ manifold of dimension at least $2$ is compact with finite fundamental group. We prove the same result for $Z^{x}$ manifolds $M$ for which all unit-speed geodesics starting at $x$ return to $x$ in uniformly bounded time. We also prove that any $Z^{x}$ manifold $(M,h)$ with $h$ analytic is a $Y^{x}_{l}$ manifold for some $l>0$. We start by defining a class of globally hyperbolic spacetimes (called observer-refocusing) such that any $Z^{x}$ manifold is the Cauchy surface of some observer-refocusing spacetime. We then prove that under suitable conditions the Cauchy surfaces of observer-refocusing spacetimes are compact with finite fundamental group and show that analytic observer-refocusing spacetimes of dimension at least $3$ are strongly refocusing. We end by stating a contact-theoretic conjecture analogous to our results in Riemannian and Lorentzian geometry.
@@ 1 Introduction
Throughout the paper, all Riemannian manifolds and Lorentzian manifolds will be assumed to be smooth, connected, and without boundary. Further all geodesics will be assumed to be maximal unless otherwise stated. If $(X,g)$ is a Lorentzian manifold, $p\in X$ and $v\in T_{p}X$, then $\alpha_{v}$ will always denote the geodesic defined through $\alpha(0)=p$ and $\alpha^{\prime}(0)=v$.
@@
Let $(M,h)$ be a complete Riemannian manifold, $x,y\in M$ and $l>0$. We say that $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold if for all geodesics $\alpha$ in $(M,h)$ with $\alpha(0)=x$ and $g(\alpha^{\prime}(0),\alpha^{\prime}(0))=1$ we have $\alpha(l)=y$. If $x=y$ we call $(M,h)$ a $Y^{x}_{l}$ manifold. Note also that every $Y^{(x,y)}_{l}$ is a $Y^{x}_{2l}$ manifold (see [15]).
@@
Let $(M,h)$ be a complete Riemannian manifold, $x,y\in M$ We say that $(M,h)$ is a $Z^{(x,y)}$ manifold if for all geodesics $\alpha$ in $(M,h)$ with $\alpha(0)=x$ there is some $t>0$ with $\alpha(t)=y$. If $x=y$ we call $(M,h)$ a $Z^{x}$ manifold. In contrast to $Y^{(x,y)}_{l}$ manifolds it is not known whether every $Z^{(x,y)}$ manifold is a $Z^{x}$ manifold.
@@
Let $(X,g)$ be a spacetime. We say that $(X,g)$ is strongly refocusing if there are $p,q\in X$ with $p\neq q$ such that all null-geodesics $\alpha$ in $(X,g)$ going through $p$ also go through $q$.
@@
Let $(X,g)$ be a spacetime, $p\in X$ and $\gamma$ some timelike curve in $X$. We say that $(X,g)$ is observer-refocusing if for all future-oriented null-geodesics $\alpha$ in $(X,g)$ with $\alpha(0)=p$ there is some $t>0$ with $\alpha(t)\in\mathrm{im}(\gamma)$.
Observer-refocusing spacetimes have the following physical interpretation. The worldline of physical observers through spacetime is a timelike curve. A spacetime is observer-refocusing, if there is some point in spacetime and some observer, such that if light is emitted from the point in spacetime in every direction, the observer will, over the course of their life, be able to see all that light. Every $Y^{(x,y)}_{l}$ manifold is a $Z^{(x,y)}$ manifold and every strongly refocusing spacetime is an observer-refocusing spacetime. (It turns out that the roles of $p$ and $q$ in Definition 1.3 are symmetric, see [19], hence we can assume without loss of generality that $q\in J^{+}(p)$ and pick $\gamma$ to be any timelike curve through $q$.)
@@
Let $(M,h)$ be a Riemannian manifold. Let $X=M\times\mathbb{R}$ and let $g=h\oplus-dt^{2}$. Then $(X,g)$ is a Lorentzian manifold. $(X,g)$ is a special case of what is more generally known as a warped product of $(M,h)$ and $(\mathbb{R},-dt^{2})$, in this case with trivial warping function. We equip $X$ with the time-orientation given by the vectorfield $\partial_{t}$ to make it a spacetime. We call $(X,g)$ the spacetime associated to the Riemannian manifold $(M,h)$.
@@
A spacetime $(X,g)$ is called causal if it admits no closed timelike curves. A spacetime is called globally hyperbolic if it is causal and if for all $p,q\in X$ the set $J^{+}(p)\cap J^{-}(q)$ is compact. A submanifold $S\subset X$ is called a Cauchy surface of $(X,g)$ if every inextendible causal curve in $X$ intersects $S$ exactly once. The existence of a Cauchy surface $S$ of a spacetime $(X,g)$ is equivalent to the spacetime being globally hyperbolic, see [18]. Further, any two Cauchy surfaces of a given spacetime $(X,g)$ are diffeomorphic and $X$ is diffeomorphic to $S\times\mathbb{R}$ for any Cauchy surface $S$, see [17].
@@
For a spacetime $(X,g)$ a smooth temporal function $\mathcal{T}:X\to\mathbb{R}$ is a smooth function with everywhere past-pointing timelike gradient. In [4], [5] Bernal and Sanchez prove that for every globally hyperbolic spacetime $X$, there are smooth, surjective temporal functions $\mathcal{T}:X\to\mathbb{R}$ such that $S_{t}\coloneqq\mathcal{T}^{-1}(t)\subset X$ is a smooth spacelike Cauchy surface for each $t\in\mathbb{R}$. Smooth temporal functions are strictly increasing along every future-directed causal curve.
@@
If $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold, then the associated spacetime $(X,g)$ is a strongly refocusing spacetime with $p=(x,0)$ and $q=(y,l)$. If $(M,h)$ is a $Z^{(x,y)}$ manifold, then the associated spacetime $(X,g)$ is an observer-refocusing spacetime with $p=(x,0)$ and $\gamma:\mathbb{R}\to X,$ $\gamma(t)=(y,t)$.
@@
The results follow immediately after realizing that, up to orientation-preserving, affine reparameterization, any future-oriented null-geodesics $\alpha$ in $(X,g)$ through $(x,0)$ is given by $\alpha(t)=(\beta(t),t)$ where $\beta$ is a unit-speed geodesic through $x$ in $(M,h)$. ∎
By proving theorems about the topology of Cauchy surfaces of globally hyperbolic observer-refocusing (resp. strongly refocusing) spacetimes, we will recover theorems about the topology of $Z^{(x,y)}$ (resp. $Y^{(x,y)}_{l}$) manifolds as a special case.
@@ 2 Main Results
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$, let $L$ be some lightsphere at $p$ (see Definition 4.29), let $\{\gamma_{i}\}_{i\in I}$ be some countable family of timelike curves and let $T>0$. If for each future-oriented lightlike geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in L$ there is some $t\in(0,T]$ and $i\in I$ such that $\alpha(t)\in\mathrm{im}(\gamma_{i})$, then $(X,g)$ has compact Cauchy surface and finite fundamental group.
2.11 Result (Proposition 4.41).
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$ which is observer-refocusing with respect to a point $p\in X$ and a timelike curve $\gamma$ such that the domain of $\gamma$ is compact, then $(X,g)$ has compact Cauchy surface and finite fundamental group.
@@
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ such that all unit-speed geodesics starting from $x$ reach $y$ in uniformly bounded time. Then $M$ is compact with finite fundamental group.
@@
Let $(X,g)$ be a globally hyperbolic, observer-refocusing spacetime of dimension at least $3$, with $g$ an analytic metric. Then $(X,g)$ is strongly refocusing.
@@
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ with $h$ an analytic metric, then $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold for some $l>0$.
@@ 3 Related results about $Y^{x}_{l}$ manifolds and refocusing spacetimes
We first review results about the topology of $Y^{x}_{l}$ manifolds and refocusing spacetimes. $Y^{x}_{l}$ manifolds, as well as $Z^{x}$ manifolds and a whole family of related notions, are discussed in [9, Definitions 7.7]. The following result is due to Bott [10] and Samelson [29].
@@
In [3] Bérard-Bergery proves a similar statement only requiring the geodesics to be closed (as opposed to periodic).
@@
Let $(M,h)$ be a $Y^{x}_{l}$ manifold of dimension at least $2$, then $M$ is compact, has finite fundamental group and the rational cohomology ring of $M$ is generated by one element.
By recent contact Bott-Samelson Theorem results of Frauenfelder, Labrousse and Schlenk [16], the conclusion of the Bérard-Bergery theorem can be strengthened to be exactly the same as the one in Bott-Samelson Theorem. Note that is it not known if every $Y^{x}_{l}$ manifold admits a metric in which all geodesics at $x$ are periodic.
@@
Let $(X,g)$ be a Lorentzian manifold. We say that $(X,g)$ is refocusing if there exists some $p\in X$ and some open set $U\ni x$ such that for all $V$ with $x\in V\subset U$ there is some $y\in X\setminus V$ such that all null-geodesics through $y$ enter $V$.
Low proved that globally hyperbolic spacetimes must have compact Cauchy surface. In [15, Theorem 11.5] Chernov and Rudyak prove that globally hyperbolic refocusing spacetimes of dimension at least $3$ have a globally hyperbolic semi-Riemannian universal cover which is also refocusing. They use this to prove the following result.
@@
Every refocusing, globally hyperbolic spacetime of dimension at least $3$ has compact Cauchy surface and finite fundamental group.
@@
3.19 Definition ($\tilde{Y}^{x}$ manifolds).
Let $(M,h)$ be a complete Riemannian manifold and let $x\in M$. We say that $M$ is a $\tilde{Y}^{x}_{l}$ manifold, if there exists some $\bar{\varepsilon}>0$ such that for all $\varepsilon$ with $0<\varepsilon<\bar{\varepsilon}$ there is some $l>\varepsilon$ and some $y$ such that for all unit-speed geodesics $\gamma$ starting at $y$ we have $d(x,\gamma(l))<\varepsilon$.
@@
Just as it is unknown whether there are $Z^{x}$ manifolds that are not $Y^{x}_{l}$ manifolds [9, Question 7.70], it is unclear whether there are $\tilde{Y}^{x}$ manifolds that are not $Y^{x}_{l}$ manifolds, whether there are refocusing spacetimes that are not strongly refocusing, and whether there are observer-refocusing spacetimes that are not strongly refocusing.
@@
Let $(X,g)$ be a spacetime and $p\in X$. We say that $(X,g)$ is very weakly refocusing if every future-pointing null-geodesic starting at $x$ eventually enters the timelike future $I^{+}(p)$.
Every $Z^{x}$ manifold is a $\tilde{Z}^{x}$ manifold but the converse is false. For example, any $n$-torus $T^{n}$ with standard metric will be a $\tilde{Z}^{x}$ manifold, but only $T^{1}$ is a $Z^{x}$ manifold. Similarly there are not observer-refocusing. We will show in Proposition 4.32 that globally hyperbolic very weakly refocusing spacetimes have compact Cauchy surface. In fact every proof of compactness of Cauchy surfaces in this paper will ultimately rely on this fact.
@@
Let $\alpha$ be a future-pointing null-geodesic with $\alpha(0)=(x,0)\eqqcolon p$. We can parametrize $\alpha$ so that $\alpha(t)=(\beta(t),t)$ for some unit-speed geodesic $\beta$ in $M$ with $\beta(0)=x$. Since $(M,h)$ is a $\tilde{Z}^{x}$ manifold, there is some $\varepsilon>0$ and some $T>\varepsilon$ such that $d(x,\beta(t))<\varepsilon$. We can assume that $d(x,\beta(T))\neq 0$ else increase $T$ by a tiny bit. Let $\tilde{\beta}$ be a unit-speed geodesic with $\tilde{\beta}(0)=x$ and $\tilde{\beta}(d(x,\beta(T))=\beta(T)$. Consider the curve $\gamma$ in $X$ defined via $\gamma(t)=(\tilde{\beta}(t),\frac{\varepsilon}{d(x,\beta(T))}t)$. It is easy to see that $\gamma$ is a timelike curve with $\gamma(0)=p$ and $\gamma(d(x,\beta(T))=(\beta(T),\varepsilon)$. Since $T>\varepsilon$ we have that $\alpha(T)=(\beta(T),T)$ is in the timelike future of $(\beta(T),\varepsilon)$, which is in the timelike future of $p=\alpha(0)=\gamma(0)$. This shows that $\alpha$ eventually enters the $I^{+}(p)$. Since this is true for all future-pointing null-geodesics $\alpha$ starting at $p$, we conclude that $(X,g)$ is very weakly refocusing. ∎
@@
In [22] Low showed that the space of future-directed, unparameterized, maximal null-geodesics $\mathcal{N}$ of a strongly causal spacetime $(X,g)$ is a smooth manifold with a natural contact structure. For globally hyperbolic manifolds, there is a natural contactomorphism from $\mathcal{N}$ to $S^{\ast}M$ for any spacelike Cauchy surface $M\subset X$ given by the following map: To each future-directed null-geodesic $\alpha$ in $X$ there is a unique point $p\in M$ and $t\in\mathbb{R}$ with $\alpha(t)=p$. Let $\pi:X=M\times\mathbb{R}\to M$ be the natural projection. Consider the vector $v\in T_{p}M$ defined through $v=d\pi(\alpha^{\prime}(t))$. Orientation preserving affine reparameterization will change $v$ by multiplication with a positive scalar. Let $[\alpha]$ denote the equivalence class of null-geodesics up to reparametrization and let $[v]$ denote the equivalence class of vectors in $TM$ up to multiplication by a positive scalar. Then the map $\mathcal{N}\to SM$, $[\alpha]\mapsto[v]$ is well-defined. It is easy to check that this map is a diffeomorphism. Since $g$ restricts to a Riemannian metric on $M$, we can identify $SM$ with $S^{\ast}M$ through $g|_{M}$. This gives us the contactomorphism $\mathcal{N}\to S^{\ast}M$. For each $p\in X$ the set $\mathrm{sky}_{p}\coloneqq\{[\alpha]\in\mathcal{N}\mid\alpha\text{ is a null-geodesic through }p\}$ is called the sky of $p$. Each sky is a Legendrian submanifold of $\mathcal{N}$, and if $M$ is any Cauchy surface through $p$ then $\mathrm{sky}_{p}$ corresponds to a fiber of $S^{\ast}M$ when identifying $\mathcal{N}\simeq S^{\ast}M$.
@@
That $(\phi_{t}(\mathrm{sky}_{p}))_{t\in\mathbb{R}}$ is a Legendrian isotopy follows from the fact that each $\phi_{t}$ is a contactomorphism. It is left to show that the isotopy is positive. Let $[\beta]\in\mathrm{sky}_{p}$. We can parametrize $\beta$ as a pre-geodesic so that $\beta(t)=(x(t),t)$. Note that $\pi(\phi_{t}([\beta]))=x(t)$. Let $\psi:U_{h_{t}}M\to S^{\ast}M$ be the natural map from the unit tangent bundle to the spherical cotangent bundle. By definition of $\phi_{t}$, there is some $\lambda>0$ with $\psi^{-1}(\phi_{t}([\beta]))=\lambda x^{\prime}(t)$. Therefore
@@
Let $X$ be a globally hyperbolic, strongly refocusing spacetimes of dimension at least $3$. Then the Cauchy surface $M$ of $X$ is compact with finite fundamental group, and its universal cover has the integral cohomology ring of a CROSS.

### v4

@@ Abstract
Let $(M,h)$ be a connected, complete Riemannian manifold, let $x\in M$ and $l>0$. Then $M$ is called a $Z^{x}$ manifold if all geodesics starting at $x$ return to $x$ and it is called a $Y^{x}_{l}$ manifold if every unit-speed geodesic starting at $x$ returns to $x$ at time $l$. It is unknown whether there are $Z^{x}$ manifolds that are not $Y^{x}_{l}$ manifolds for any $l>0$. By the Bérard-Bergery theorem, any $Y^{x}_{l}$ manifold of dimension at least $2$ is compact with finite fundamental group. We prove the same result for $Z^{x}$ manifolds $M$ for which all unit-speed geodesics starting at $x$ return to $x$ in uniformly bounded time. We also prove that any $Z^{x}$ manifold $(M,h)$ with $h$ analytic is a $Y^{x}_{l}$ manifold for some $l>0$. We start by defining a class of globally hyperbolic spacetimes (called observer-refocusing) such that any $Z^{x}$ manifold is the Cauchy surface of some observer-refocusing spacetime. We then prove that under suitable conditions the Cauchy surfaces of observer-refocusing spacetimes are compact with finite fundamental group and show that analytic observer-refocusing spacetimes of dimension at least $3$ are strongly refocusing. We end by stating a contact-theoretic conjecture analogous to our results in Riemannian and Lorentzian geometry.
@@ 1 Introduction
Throughout the paper, all Riemannian manifolds and Lorentzian manifolds will be assumed to be smooth, connected, and without boundary. Further, all geodesics will be assumed to be maximal unless otherwise stated. If $(X,g)$ is a Lorentzian manifold, $p\in X$ and $v\in T_{p}X$, then $\alpha_{v}$ will always denote the geodesic defined through $\alpha_{v}(0)=p$ and $\alpha_{v}^{\prime}(0)=v$.
@@
Let $(M,h)$ be a complete Riemannian manifold, $x,y\in M$ and $l>0$. We say that $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold if for all geodesics $\alpha$ in $(M,h)$ with $\alpha(0)=x$ and $h(\alpha^{\prime}(0),\alpha^{\prime}(0))=1$ we have $\alpha(l)=y$. If $x=y$ we call $(M,h)$ a $Y^{x}_{l}$ manifold. Note also that every $Y^{(x,y)}_{l}$ manifold is a $Y^{x}_{2l}$ manifold (see [16]).
@@
Let $(M,h)$ be a complete Riemannian manifold, $x,y\in M$. We say that $(M,h)$ is a $Z^{(x,y)}$ manifold if for all geodesics $\alpha$ in $(M,h)$ with $\alpha(0)=x$ there is some $t>0$ with $\alpha(t)=y$. If $x=y$ we call $(M,h)$ a $Z^{x}$ manifold. In contrast to $Y^{(x,y)}_{l}$ manifolds, it is not known whether every $Z^{(x,y)}$ manifold is a $Z^{x}$ manifold.
@@
Let $(X,g)$ be a spacetime and let $p,q\in X$ with $p\neq q$. We say that $(X,g)$ is strongly refocusing with respect to $p$ and $q$ if all null-geodesics $\alpha$ in $(X,g)$ going through $p$ also go through $q$. We say that a spacetime $(X,g)$ is strongly refocusing if there are $p\neq q$ such that $(X,g)$ is strongly refocusing with respect to $p$ and $q$.
@@
Let $(X,g)$ be a spacetime, $p\in X$ and $\gamma$ some timelike curve in $X$. We say that $(X,g)$ is future (resp. past) observer-refocusing with respect to $p$ and $\gamma$ if for all future-oriented (resp. past-oriented) null-geodesics $\alpha$ in $(X,g)$ with $\alpha(0)=p$ there is some $t>0$ with $\alpha(t)\in\mathrm{im}(\gamma)$. We say a spacetime is future (resp. past) observer-refocusing if it is future (resp. past) observer-refocusing with respect to some point $p\in X$ and some timelike curve $\gamma$ in $X$.
Future observer-refocusing spacetimes have the following physical interpretation. The worldline of physical observers through spacetime is a future-directed timelike curve. A spacetime is observer-refocusing, if there is some point in spacetime and some observer, such that if light is emitted from the point in spacetime in every direction, the observer will, over the course of their life, be able to see all that light. Every $Y^{(x,y)}_{l}$ manifold is a $Z^{(x,y)}$ manifold and every strongly refocusing spacetime is a future observer-refocusing spacetime and a past observer-refocusing spacetime. It turns out that the roles of $p$ and $q$ in Definition 1.3 are symmetric, see [21]. Without loss of generality, we may assume that $q\in J^{+}(p)$. By picking $\gamma$ to be any timelike curve through $q$, the spacetime will be future observer-refocusing with respect to $p$ and $\gamma$. Conversely, if we pick $\gamma$ to be any timelike curve through $p$, then the spacetime will be past observer-refocusing with respect to $q$ and $\gamma$. Throughout the rest of the paper we will use ”observer-refocusing” to mean future observer-refocusing. Analogous results for past observer-refocusing spacetimes can be obtained by reversing the time-orientation of the spacetime.
@@
Let $(M,h)$ be a Riemannian manifold. Let $X=M\times\mathbb{R}$ and let $g=h\oplus-dt^{2}$. Then $(X,g)$ is a Lorentzian manifold. We equip $X$ with the time-orientation given by the vectorfield $\partial_{t}$ to make it a spacetime. We call $(X,g)$ the spacetime associated to the Riemannian manifold $(M,h)$.
@@
A spacetime $(X,g)$ is called causal if it admits no closed timelike curves. A spacetime is called globally hyperbolic if it is causal and if for all $p,q\in X$ the set $J^{+}(p)\cap J^{-}(q)$ is compact. A submanifold $S\subset X$ is called a Cauchy surface of $(X,g)$ if every inextendible causal curve in $X$ intersects $S$ exactly once. The existence of a Cauchy surface $S$ of a spacetime $(X,g)$ is equivalent to the spacetime being globally hyperbolic, see [19], [20]. Further, any two Cauchy surfaces of a given spacetime $(X,g)$ are diffeomorphic and $X$ is diffeomorphic to $S\times\mathbb{R}$ for any Cauchy surface $S$, see [19], [5].
@@
For a spacetime $(X,g)$ a smooth temporal function $\mathcal{T}:X\to\mathbb{R}$ is a smooth function with everywhere past-pointing timelike gradient. In [5], [6] Bernal and Sánchez prove that for every globally hyperbolic spacetime $X$, there are smooth, surjective temporal functions $\mathcal{T}:X\to\mathbb{R}$ such that $S_{t}\coloneqq\mathcal{T}^{-1}(t)\subset X$ is a smooth spacelike Cauchy surface for each $t\in\mathbb{R}$. Smooth temporal functions are strictly increasing along every future-directed causal curve.
@@
If $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold, then the associated spacetime $(X,g)$ is a strongly refocusing spacetime with $p=(x,0)$ and $q=(y,l)$. If $(M,h)$ is a $Z^{(x,y)}$ manifold, then the associated spacetime $(X,g)$ is an observer-refocusing spacetime with $p=(x,0)$ and observer $\gamma:\mathbb{R}\to X,$ $\gamma(t)=(y,t)$.
@@
The results follow immediately after realizing that, up to orientation-preserving, affine reparameterization, any future-oriented null-geodesic $\alpha$ in $(X,g)$ through $(x,0)$ is given by $\alpha(t)=(\beta(t),t)$ where $\beta$ is a unit-speed geodesic through $x$ in $(M,h)$. ∎
[…]
@@ 2 Main Results
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$, let $L$ be some lightsphere at $p$ (see Definition 4.29), let $\{\gamma_{i}\}_{i\in I}$ be some countable family of timelike curves and let $T>0$. If for each future-oriented null geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in L$ there is some $t\in(0,T]$ and $i\in I$ such that $\alpha(t)\in\mathrm{im}(\gamma_{i})$, then $(X,g)$ has a compact Cauchy surface and finite fundamental group.
@@
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$ which is observer-refocusing with respect to a point $p\in X$ and a timelike curve $\gamma$ such that the domain of $\gamma$ is compact. Then $(X,g)$ has a compact Cauchy surface and finite fundamental group.
2.12 Result (Corollary 4.44).
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ such that all unit-speed geodesics starting at $x$ reach $y$ in uniformly bounded time. Then $M$ is compact with finite fundamental group.
@@
Let $(X,g)$ be a globally hyperbolic, observer-refocusing spacetime of dimension at least $3$ with an analytic metric $g$. Then $(X,g)$ is strongly refocusing.
@@
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ with an analytic metric $h$, then $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold for some $l>0$.
@@ 3 Related results about $Y^{x}_{l}$ manifolds and refocusing spacetimes
We first review results about the topology of $Y^{x}_{l}$ manifolds and refocusing spacetimes. $Y^{x}_{l}$ manifolds, as well as $Z^{x}$ manifolds and a whole family of related notions, are discussed in [10, Definitions 7.7]. The following result is due to Bott [11] and Samelson [32].
@@
In [4] Bérard-Bergery proves a similar statement only requiring the geodesics to revisit $x$ after a common time $l$, removing the requirement that the geodesics return with the same velocity they started with.
@@
Let $(M,h)$ be a $Y^{x}_{l}$ manifold of dimension at least $2$. Then $M$ is compact, has finite fundamental group and the rational cohomology ring of $M$ is generated by one element.
By recent contact Bott-Samelson Theorem results of Frauenfelder, Labrousse and Schlenk [18], the conclusion of the Bérard-Bergery theorem can be strengthened to be exactly the same as the one in the Bott-Samelson Theorem.
@@
Let $(X,g)$ be a Lorentzian manifold. We say that $(X,g)$ is refocusing if there exists some $p\in X$ and some open set $U\ni p$ such that for all $V$ with $p\in V\subset U$ there is some $q\in X\setminus V$ such that all null-geodesics through $q$ enter $V$.
Low proved that refocusing globally hyperbolic spacetimes must have a compact Cauchy surface. In [16, Theorem 11.5] Chernov and Rudyak prove that globally hyperbolic refocusing spacetimes of dimension at least $3$ have a globally hyperbolic semi-Riemannian universal cover which is also refocusing. They use this to prove the following result.
@@
Every refocusing, globally hyperbolic spacetime of dimension at least $3$ has a compact Cauchy surface and finite fundamental group.
@@
3.19 Definition ($\tilde{Y}^{x}_{l}$ manifolds).
Let $(M,h)$ be a complete Riemannian manifold and let $x\in M$. We say that $M$ is a $\tilde{Y}^{x}_{l}$ manifold if there exists some $\bar{\varepsilon}>0$ such that for all $\varepsilon$ with $0<\varepsilon<\bar{\varepsilon}$ there is some $l>\varepsilon$ and some $y$ such that for all unit-speed geodesics $\gamma$ starting at $y$ we have $d(x,\gamma(l))<\varepsilon$.
@@
Just as it is unknown whether there are $Z^{x}$ manifolds that are not $Y^{x}_{l}$ manifolds [10, Question 7.70], it is unclear whether there are $\tilde{Y}^{x}$ manifolds that are not $Y^{x}_{l}$ manifolds, whether there are refocusing spacetimes that are not strongly refocusing, and whether there are observer-refocusing spacetimes that are not strongly refocusing.
@@
Let $(X,g)$ be a spacetime and $p\in X$. We say that $(X,g)$ is very weakly refocusing if every future-pointing null-geodesic starting at $p$ eventually enters the timelike future $I^{+}(p)$.
Every $Z^{x}$ manifold is a $\tilde{Z}^{x}$ manifold but the converse is false. For example, any $n$-torus $T^{n}$ with standard metric will be a $\tilde{Z}^{x}$ manifold, but only $T^{1}$ is a $Z^{x}$ manifold. We will show in Proposition 4.32 that globally hyperbolic very weakly refocusing spacetimes have a compact Cauchy surface. In fact every proof of compactness of Cauchy surfaces in this paper will ultimately rely on this fact.
@@
Let $\alpha$ be a future-pointing null-geodesic with $\alpha(0)=(x,0)\eqqcolon p$. We can parametrize $\alpha$ so that $\alpha(t)=(\beta(t),t)$ for some unit-speed geodesic $\beta$ in $M$ with $\beta(0)=x$. Since $(M,h)$ is a $\tilde{Z}^{x}$ manifold, there is some $\varepsilon>0$ and some $T>\varepsilon$ such that $d(x,\beta(T))<\varepsilon$. We can assume that $d(x,\beta(T))\neq 0$ else increase $T$ by a tiny bit. Let $\tilde{\beta}$ be a unit-speed geodesic with $\tilde{\beta}(0)=x$ and $\tilde{\beta}(d(x,\beta(T)))=\beta(T)$. Consider the curve $\gamma$ in $X$ defined via $\gamma(t)=(\tilde{\beta}(t),\frac{\varepsilon}{d(x,\beta(T))}t)$. It is easy to see that $\gamma$ is a timelike curve with $\gamma(0)=p$ and $\gamma(d(x,\beta(T)))=(\beta(T),\varepsilon)$. Since $T>\varepsilon$ we have that $\alpha(T)=(\beta(T),T)$ is in the timelike future of $(\beta(T),\varepsilon)$, which is in the timelike future of $p=\alpha(0)=\gamma(0)$. This shows that $\alpha$ eventually enters the $I^{+}(p)$. Since this is true for all future-pointing null-geodesics $\alpha$ starting at $p$, we conclude that $(X,g)$ is very weakly refocusing. ∎
@@
In [24] Low showed that the space of future-directed, unparameterized, maximal null-geodesics $\mathcal{N}$ of a strongly causal spacetime $(X,g)$ is a smooth manifold with a natural contact structure. For globally hyperbolic manifolds, there is a natural contactomorphism from $\mathcal{N}$ to $S^{\ast}M$ for any spacelike Cauchy surface $M\subset X$ given by the following map: To each future-directed null-geodesic $\alpha$ in $X$ there is a unique point $p\in M$ and $t\in\mathbb{R}$ with $\alpha(t)=p$. Let $\pi:X=M\times\mathbb{R}\to M$ be the natural projection. Consider the vector $v\in T_{p}M$ defined through $v=d\pi(\alpha^{\prime}(t))$. Orientation-preserving affine reparameterization will change $v$ by multiplication with a positive scalar. Let $\alpha]$ denote the equivalence class of null-geodesics up to reparametrization and let $[v]$ denote the equivalence class of vectors in $TM$ up to multiplication by a positive scalar. Then the map $\mathcal{N}\to SM$, $\alpha]\mapsto[v]$ is well-defined. It is easy to check that this map is a diffeomorphism. Since $g$ restricts to a Riemannian metric on $M$, we can identify $SM$ with $S^{\ast}M$ through $g|_{M}$. This gives us the contactomorphism $\mathcal{N}\to S^{\ast}M$. For each $p\in X$ the set $\mathrm{sky}_{p}\coloneqq\{[\alpha]\in\mathcal{N}\mid\alpha\text{ is a null-geodesic through }p\}$ is called the sky of $p$. Each sky is a Legendrian submanifold of $\mathcal{N}$, and if $M$ is any Cauchy surface through $p$ then $\mathrm{sky}_{p}$ corresponds to a fiber of $S^{\ast}M$ when identifying $\mathcal{N}\simeq S^{\ast}M$.
@@
That $(\phi_{t}(\mathrm{sky}_{p}))_{t\in\mathbb{R}}$ is a Legendrian isotopy follows from the fact that each $\phi_{t}$ is a contactomorphism. It is left to show that the isotopy is positive. Let $\beta]\in\mathrm{sky}_{p}$. We can parametrize $\beta$ as a pre-geodesic so that $\beta(t)=(x(t),t)$. Note that $\pi(\phi_{t}([\beta]))=x(t)$. Let $\psi:U_{h_{t}}M\to S^{\ast}M$ be the natural map from the unit tangent bundle to the spherical cotangent bundle. By definition of $\phi_{t}$, there is some $\lambda>0$ with $\psi^{-1}(\phi_{t}([\beta]))=\lambda x^{\prime}(t)$. Therefore
@@
Let $X$ be a globally hyperbolic, strongly refocusing spacetime of dimension at least $3$. Then the Cauchy surface $M$ of $X$ is compact with finite fundamental group, and its universal cover has the integral cohomology ring of a CROSS.